OpenStudy (anonymous):

check the rolle's theorem and the mean value theorem and find a value of C that makes the appropriate conclusion true. F(x)=sin(x) [-pi, 0]

OpenStudy (anonymous):

it must satisfy it, because it is differentiable and continuous and also \(\sin(-\pi)=\sin(0)\)

OpenStudy (anonymous):

but how do you go on finding "C"?